OpenStudy (anonymous):

In a cafeteria, the ratio of the number of girls to the number of boys was 6:7. When 117 girls entered, the new ratio of the number of girls to boys was 7:6. How many pupils were in the auditorium at first.

OpenStudy (anonymous):

100. i think.

OpenStudy (anonymous):

217, you have to add 117 amount of girls plus the amount of boys. You get 100 from 117/7 then times 6

OpenStudy (anonymous):

so its 217?

OpenStudy (anonymous):

Should be yes.

OpenStudy (anonymous):

But how did you get that..?

OpenStudy (anonymous):

You take 217/7 and then times it by 6 to get how many guys there were then add that number to the 117 girls already in there since it says amount of pupils

OpenStudy (anonymous):

217/7 = 31 * 6 = 186 + 117 = 303

OpenStudy (anonymous):

So sorry, I meant 117/7 * 6= 100

OpenStudy (anonymous):

so its 217... hmm

OpenStudy (anonymous):

In the way I interpret the question yes that is correct.

OpenStudy (anonymous):

it was asking how many were in the auditorium "AT FIRST"...

OpenStudy (anonymous):

Yes and if you look at it the ratio went from 7:6 to 6:7, therefore the number of people should stay the same. I would think.

OpenStudy (anonymous):

it went from 6:7 to 7:6

OpenStudy (anonymous):

Eh, same different. The ratio stayed the same in the long run. If it were to go from 6:7 to 3:4 then I would think there is more to it.

OpenStudy (anonymous):

Ok so its still 217...

OpenStudy (anonymous):

I believe so. Ask around some more if you'd like a second opinion.

OpenStudy (anonymous):

No i think your rig ht because that makes sense and your the only one replying